# bayesOptimalControl.bib

### v1

@@ Appendix A Details on the characteristics
As discussed in Section 5.1, it is necessary to truncate the originally unbounded domain to obtain a numerical solution for the proposed PDE (5.1a). To effectively truncate the domain and construct a convergent numerical scheme, it is crucial to understand the propagation of information within the PDE. To achieve this, we compute the characteristics of equation (5.1a) by following the methodology outlined in Section 3.2 of [evans@2010]. These characteristics provide the curves along which information flows from initial points throughout the domain in first-order PDEs. By knowing these curves, we can appropriately select a domain where the characteristic lines flow from inside the domain to the outside, ensuring that the numerical scheme is well-posed and convergent.
@@
(A.2) $\begin{cases}z(\tau)=\frac{b^{2}}{2\theta}+C_{1}(m_{0},z_{0})e^{2\theta\tau},\\ q(\tau)=\frac{1}{2\theta}+C_{2}(m_{0},z_{0})e^{-2\theta\tau},\\ p(\tau)=C_{3}(m_{0},z_{0})e^{\sqrt{\theta^{2}+1/C}\tau}+C_{4}(m_{0},z_{0})e^{-\sqrt{\theta^{2}+1/C}\tau},\\ m(\tau)=(\theta+\sqrt{\theta^{2}+1/C})\frac{C_{3}(m_{0},z_{0})}{2}e^{\sqrt{\theta^{2}+1/C}\tau}+(\theta-\sqrt{\theta^{2}+1/C})\frac{C_{4}(m_{0},z_{0})}{2}e^{-\sqrt{\theta^{2}+1/C}\tau}\end{cases}$
@@ Appendix B Derivation of the monotonicity condition
In Section 5.2 we construct the numerical scheme which is guaranteed to converge to the true solution, provided that it is consistent and monotone. In this Appendix we derive the condition which ensures that our scheme is monotone. To do that, we check directly under which conditions the function $G$ is monotone in each of its arguments $U^{n}_{i-1,j}$, $U^{n}_{i,j}$, $U^{n}_{i+1,j}$, $U^{n}_{i,j-1}$, $U^{n}_{i,j+1}$.
@@
Since $U^{n+1}_{i,j}$ is an increasing function, so $\frac{\partial U^{n+1}_{i,j}}{\partial U^{n}_{i,j}}\geq 0$. Recalling that $\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha)(\alpha-\alpha_{0})\geq 0$ and expanding on $h_{1}(m_{i},\alpha,\beta)$ we get
(B.8) $h_{1}(m_{i},z_{j},\alpha,\beta)$ =$=\begin{cases}\frac{1}{\Delta m}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\beta),~{}~{}~{}~{}\text{if}~{}\beta,\alpha\leq\alpha_{0},\\ \frac{1}{\Delta m}\Big{(}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\beta)-\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha)\Big{)},~{}~{}~{}~{}\text{if}~{}\beta\leq\alpha_{0},~{}\alpha\geq\alpha_{0},\\ 0,~{}~{}~{}~{}\text{if}~{}\beta\geq\alpha_{0},~{}\alpha\leq\alpha_{0},\\ \frac{1}{\Delta m}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha),~{}~{}~{}~{}\text{if}~{}\beta,\alpha\geq\alpha_{0}\end{cases}$
(B.9) =$=\begin{cases}\frac{1}{\Delta m}\Big{|}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\beta)\Big{|},~{}~{}~{}~{}\text{if}~{}\beta,\alpha\leq\alpha_{0},\\ \frac{1}{\Delta m}\Big{(}\Big{|}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\beta)\Big{|}+\Big{|}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha)\Big{|}\Big{)},~{}~{}~{}~{}\text{if}~{}\beta\leq\alpha_{0},~{}\alpha\geq\alpha_{0},\\ 0,~{}~{}~{}~{}\text{if}~{}\beta\geq\alpha_{0},~{}\alpha\leq\alpha_{0},\\ \frac{1}{\Delta m}\Big{|}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha)\Big{|},~{}~{}~{}~{}\text{if}~{}\beta,\alpha\geq\alpha_{0}\end{cases}$
@@
On a domain defined by $[m_{0},m_{I}]\times[z_{0},z_{J}]$ with $m_{0}<0$, we have the following bounds

### v2

@@ Appendix A Details on the characteristics
As discussed in Section 5.1.1, it is necessary to truncate the originally unbounded domain to obtain a numerical solution for the proposed PDE (5.38a). To effectively truncate the domain and construct a convergent numerical scheme, it is crucial to understand the propagation of information within the PDE. To achieve this, we compute the characteristics of equation (5.38a) by following the methodology outlined in Section 3.2 of [Eva10]. These characteristics provide the curves along which information flows from initial points throughout the domain in first-order PDEs. By knowing these curves, we can appropriately select a domain where the characteristic lines flow from inside the domain to the outside, ensuring that the numerical scheme is well-posed and convergent.
@@
(A.2) $\begin{cases}z(\tau)=\frac{b^{2}}{2\theta}+C_{1}(m_{0},z_{0})e^{2\theta\tau},\\ q(\tau)=\frac{1}{2\theta}+C_{2}(m_{0},z_{0})e^{-2\theta\tau},\\ p(\tau)=C_{3}(m_{0},z_{0})e^{\sqrt{\theta^{2}+1/C}\tau}+C_{4}(m_{0},z_{0})e^{-\sqrt{\theta^{2}+1/C}\tau},\\ m(\tau)=(\theta+\sqrt{\theta^{2}+1/C})\frac{C_{3}(m_{0},z_{0})}{2}e^{\sqrt{\theta^{2}+1/C}\tau}+...\\ \quad\quad\quad+(\theta-\sqrt{\theta^{2}+1/C})\frac{C_{4}(m_{0},z_{0})}{2}e^{-\sqrt{\theta^{2}+1/C}\tau}\end{cases}$
@@ Appendix B Derivation of the monotonicity condition
In Section 5.1.2 we construct the numerical scheme which is guaranteed to converge to the true solution, provided that it is consistent and monotone. In this Appendix we derive the condition which ensures that our scheme is monotone. To do that, we check directly under which conditions the function $G$ is monotone in each of its arguments $U^{n}_{i-1,j}$, $U^{n}_{i,j}$, $U^{n}_{i+1,j}$, $U^{n}_{i,j-1}$, $U^{n}_{i,j+1}$.
@@
Since $U^{n+1}_{i,j}$ is an increasing function, so $\frac{\partial U^{n+1}_{i,j}}{\partial U^{n}_{i,j}}\geq 0$. Recalling that $\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha)(\alpha-\alpha_{0})\geq 0$ and expanding on $h_{1}(m_{i},z_{j},\alpha,\beta)$ we get
(B.8) $h_{1}(m_{i},z_{j},\alpha,\beta)$ $=\begin{cases}-\frac{1}{\Delta m}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\beta),~~~~\text{if}~\beta,\alpha\leq\alpha_{0},\\ -\frac{1}{\Delta m}\Big{(}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\beta)-\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha)\Big{)},~~~~\text{if}~\beta\leq\alpha_{0},~\alpha\geq\alpha_{0},\\ 0,~~~~\text{if}~\beta\geq\alpha_{0},~\alpha\leq\alpha_{0},\\ \frac{1}{\Delta m}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha),~~~~\text{if}~\beta,\alpha\geq\alpha_{0}\end{cases}$
(B.9) $=\begin{cases}\frac{1}{\Delta m}\Big{|}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\beta)\Big{|},~~~~\text{if}~\beta,\alpha\leq\alpha_{0},\\ \frac{1}{\Delta m}\Big{(}\Big{|}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\beta)\Big{|}+\Big{|}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha)\Big{|}\Big{)},~~~~\text{if}~\beta\leq\alpha_{0},~\alpha\geq\alpha_{0},\\ 0,~~~~\text{if}~\beta\geq\alpha_{0},~\alpha\leq\alpha_{0},\\ \frac{1}{\Delta m}\Big{|}\frac{\partial H_{1}}{\partial(\partial\hat{U}/\partial m)}(m_{i},z_{j},\alpha)\Big{|},~~~~\text{if}~\beta,\alpha\geq\alpha_{0}\end{cases}$
@@
On a domain defined by $m_{0},m_{I}]\times[z_{0},z_{J}]$ with $m_{0}<0$, we have the following bounds